Polaski Company manufactures and sells a single product called a Ret. Operating at capacity, the company can produce and sell 34,000 Rets per year. Costs associated with this level of production and sales are given below:
Units Total
Direct materials 15
Direct labor 8
Variable manufacturing overhead 2
Fixed manufacturing overhead 9
Variable selling expense 1
Fixed selling expense 4
Total cost 39
The Rets normally sell for $56 each. Fixed manufacturing overhead is constant at $306,000 per year within the range of 29,000 through 34,000 Rets per year.

Requirement 1:

Assume that due to a recession, Polaski Company expects to sell only 29,000 Rets through regular channels next year. A large retail chain has offered to purchase 3,100 Rets if Polaski is willing to accept a 11% discount off the regular price. There would be no sales commissions on this order; thus, variable selling expenses would be slashed by 74%. However, Polaski Company would have to purchase a special machine to engrave the retail chain's name on the 3,100 units. This machine would cost $3,100. Polaski Company has no assurance that the retail chain will purchase additional units in the future. Calculate the net increase/decrease in profits next year if this special order is accepted. (Input the amount as positive value. Omit the "$" sign in your response.)

Requirement 2:

Assume again that Polaski Company expects to sell only 29,000 Rets through regular channels next year. The U.S. Army would like to make a one-time-only purchase of 3,100 Rets. The Army would pay a fixed fee of $1.77 per Ret, and it would reimburse Polaski Company for all costs of production (variable and fixed) associated with the units. Because the army would pick up the Rets with its own trucks, there would be no variable selling expenses associated with this order. If Polaski Company accepts the order, by how much will profits increase or decrease for the year? (Input the amount as positive value. Omit the "$" sign in your response.)

Requirement 3:

Assume the same situation as that described in Requirement (2) above, except that the company expects to sell 34,000 Rets through regular channels next year. Thus, accepting the U.S. Army's order would require giving up regular sales of 3,100 Rets. If the Army's order is accepted, by how much will profits increase or decrease from what they would be if the 3,100 Rets were sold through regular channels? (Input the amount as positive value. Omit the "$" sign in your response.)
Answer #1
Concepts and reason

Special Order Decision: It refers to a decision as whether to accept or reject the special order at a price below the usual market price. It uses incremental analysis to compare the special order price with the relevant costs to check whether the profit is generated.

Incremental analysis: It is a type of decision making technique used by the company in determining the effect of net income and cost in choosing between various alternatives.

Fundamentals

Sales Revenue: Sales Revenue is the revenue generated by the company when products are sold in the market.

SalesRevenue=Numberofunits×Sellingprice{\rm{Sales Revenue}} = {\rm{Number of units }} \times {\rm{ Selling price}}

Variable cost: These costs vary with the number of units produced or for the services provided. For example, the labor costs increase if the number of labor hours is increased, and the labor costs decrease if the number of labor hours is decreased. E.g.: Material cost, Labor cost, variable overhead.

Total material cost: It is the total material cost incurred by the company in manufacturing total units in a particular accounting year.

DirectMaterialsCost=Numberofunits×Directmaterialsperunit{\rm{Direct Materials Cost = Number of units }} \times {\rm{ Direct materials per unit}}

Total units manufactured: It is the total number of units that are manufactured by the company in a particular accounting year.

Special order units: It is the number of units that the company has accepted to deliver in a particular period of time.

Total labor cost: It is the total labor cost incurred by the company in manufacturing total units by total labor in a particular accounting year.

Total variable overhead: It is the total variable overhead incurred by the company in manufacturing total units by total labor in a particular accounting year.

Fixed cost: These are the costs which remain constant throughout the process of manufacturing or for the services rendered. They are incurred irrespective of number of units produced. For example, factory rent. The rent should be paid, if the factory produces 100 units or 1,000 units or if no units are produced.

Net income: The net income is the excess of revenues over expenses after adjusting for depreciation and taxes. Net income will appear in the bottom of the income statement of the company.

The formula to calculate the variable manufacturing overhead is as follows:

Variablemanufacturingoverhead}=[Numberofunits×Variablemanufacturingoverheadperunit]\left. \begin{array}{l}\\{\rm{Variable manufacturing }}\\\\{\rm{overhead }}\\\end{array} \right\}{\rm{ = }}\left[ \begin{array}{l}\\{\rm{Number of units }}\\\\ \times {\rm{ Variable manufacturing overhead per unit}}\\\end{array} \right]

The formula to calculate the variable selling expenses is as follows:

Variablesellingexpense}=[Numberofunits×Variablesellingexpenseperunit]\left. \begin{array}{l}\\{\rm{Variable }}\\\\{\rm{selling expense }}\\\end{array} \right\}{\rm{ = }}\left[ \begin{array}{l}\\{\rm{Number of units }}\\\\ \times {\rm{ Variable selling expense per unit}}\\\end{array} \right]

(1)

The Statement of Net income of P Company is calculated as follows:

Regular Sales |Special Order
3,100 Units
$154,504
Particulars
29,000 units
Sales (1)
Less: Variable costs
$1,624,000
Direct M

Therefore, the increase in net income under special order is calculated as 73,098.

(Working notes :)

The Sales for regular channel is calculated as under:

Sales(RegularChannel)=Numberofunits×Sellingprice=29,000×$56=$16,24,000\begin{array}{c}\\{\rm{Sales }}\left( {{\rm{Regular Channel}}} \right) = {\rm{Number of units }} \times {\rm{ Selling price}}\\\\{\rm{ = 29,000}}\,\, \times \,\,\$ 56\\\\ = \,\,\$ 16,24,000\\\end{array}

The Sales for special order is calculated as under:

Sales(Specialorder)=Specialorderunits×Sellingprice=3,100×[$56×(111%)]=3,100×[$56×89%]=$154,504\begin{array}{c}\\{\rm{Sales }}\left( {{\rm{Special order}}} \right) = {\rm{Special order units }} \times {\rm{ Selling price}}\\\\{\rm{ = 3,100}}\,\, \times \,\,\left[ {\$ 56 \times \,\,\left( {\,1--\,11\% } \right)\,} \right]\\\\ = \,\,{\rm{3,100}}\,\, \times \,\,\left[ {\$ 56\,\, \times \,\,89\% \,} \right]\,\\\\ = \,\,\$ 154,504\\\end{array}

The direct materials for regular sales are calculated as follows:

DirectMaterials(Regularsales)=Numberofunits×Directmaterialsperunit=29,000×$15=$435,000\begin{array}{c}\\{\rm{Direct Materials }}\left( {{\rm{Regular sales}}} \right){\rm{ = Number of units }} \times {\rm{ Direct materials per unit}}\\\\{\rm{ = 29,000}} \times \,{\rm{\$ 15}}\\\\{\rm{ = }}\,\,{\rm{\$ 435,000}}\\\end{array}

The direct materials for special order are calculated as follows:

DirectMaterials(Specialorder)=Specialorderunits×Directmaterialsperunit=3,100×$15=$46,500\begin{array}{c}\\{\rm{Direct Materials }}\left( {{\rm{Special order}}} \right){\rm{ = Special order units }} \times {\rm{ Direct materials per unit}}\\\\{\rm{ = 3,100}} \times \,{\rm{\$ 15}}\\\\{\rm{ = }}\,\,{\rm{\$ 46,500}}\\\end{array}

The direct labor for regular sales is calculated as follows:

DirectLabor(Regularsales)=Numberofunits×Directlaborperunit=29,000×$8=$232,000\begin{array}{c}\\{\rm{Direct Labor }}\left( {{\rm{Regular sales}}} \right){\rm{ = Number of units }} \times {\rm{ Direct labor per unit}}\\\\{\rm{ = 29,000}} \times \,\$ 8\\\\{\rm{ = }}\,\,\$ 232{\rm{,000}}\\\end{array}

The direct labor for special order is calculated as follows:

DirectLabor(Specialorder)=Specialorderunits×Directlaborperunit=3,100×$8=$24,800\begin{array}{c}\\{\rm{Direct Labor }}\left( {{\rm{Special order}}} \right){\rm{ = Special order units }} \times {\rm{ Direct labor per unit}}\\\\{\rm{ = 3,100}} \times \,\$ 8\\\\{\rm{ = }}\,\,\$ 24,800\\\end{array}

The variable manufacturing overhead for regular sales is calculated as follows:

Variablemanufacturingoverhead(Regularsales)}=[Numberofunits×Variablemanufacturingoverheadperunit]=29,000×$2=$58,000\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Variable manufacturing }}\\\\{\rm{overhead }}\left( {{\rm{Regular sales}}} \right)\\\end{array} \right\}{\rm{ = }}\left[ \begin{array}{l}\\{\rm{Number of units }}\\\\ \times {\rm{ Variable manufacturing overhead per unit}}\\\end{array} \right]\\\\{\rm{ = 29,000}}\,\, \times \,\,{\rm{\$ 2}}\\\\{\rm{ = }}\,\,{\rm{\$ 58,000}}\\\end{array}

The variable manufacturing overhead for special order is calculated as follows:

Variablemanufacturingoverhead(Specialorder)}=[Specialorderunits×Variablemanufacturingoverheadperunit]=3,100×$2=$6,200\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Variable manufacturing }}\\\\{\rm{overhead }}\left( {{\rm{Special order}}} \right)\\\end{array} \right\}{\rm{ = }}\left[ \begin{array}{l}\\{\rm{Special order units }}\\\\ \times {\rm{ Variable manufacturing overhead per unit}}\\\end{array} \right]\\\\{\rm{ = 3,100}}\,\, \times \,\,{\rm{\$ 2}}\\\\{\rm{ = }}\,\,{\rm{\$ 6,200}}\\\end{array}

The variable selling expenses for regular sales is calculated as follows:

Variablesellingexpense(Regularsales)}=[Numberofunits×Variablesellingexpenseperunit]=29,000×$1=$29,000\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Variable }}\\\\{\rm{selling expense }}\left( {{\rm{Regular sales}}} \right)\\\end{array} \right\}{\rm{ = }}\left[ \begin{array}{l}\\{\rm{Number of units }}\\\\ \times {\rm{ Variable selling expense per unit}}\\\end{array} \right]\\\\{\rm{ = 29,000}}\,\, \times \,\,{\rm{\$ 1}}\\\\{\rm{ = }}\,\,{\rm{\$ 29,000}}\\\end{array}

The variable selling expenses for special order is calculated as follows:

Variablesellingexpense(Specialorder)}=[Specialorderunits×Variablesellingexpenseperunit]=3,100×$1×[(174%)]=3,100×$1×26%=$806\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Variable }}\\\\{\rm{selling expense }}\left( {{\rm{Special order}}} \right)\\\end{array} \right\}{\rm{ = }}\left[ \begin{array}{l}\\{\rm{Special order units }}\\\\ \times {\rm{ Variable selling expense per unit}}\\\end{array} \right]\\\\{\rm{ = 3,100}}\,\, \times \,\,{\rm{\$ 1}}\,\, \times \,\left[ {\left( {1--74\% } \right)} \right]\\\\ = \,{\rm{3,100}}\,\, \times \,\,{\rm{\$ 1}}\,\, \times 26\% \\\\{\rm{ = }}\,\,{\rm{\$ 806}}\\\end{array}

(2)

The Contribution gained due to special order is calculated as follows:

Contribution(Specialorder)=Specialorderunits×[Reimbursementprice+Fixedfee]=3,100units×$[9+1.77]=3,100×$10.77=$33,387\begin{array}{c}\\{\rm{Contribution }}\left( {S{\rm{pecial order}}} \right) = \,\,Special{\rm{ order units }} \times {\rm{ }}\left[ {{\rm{Reimbursement price + Fixed fee}}} \right]\\\\ = \,\,3,100\,units{\rm{ }} \times {\rm{ \$ }}\left[ {9\,{\rm{ + }}\,{\rm{1}}{\rm{.77}}\,} \right]\\\\ = \,\,3,100\,\, \times \,\$ 10.77\\\\ = \,\,\$ 33,387\\\end{array}

Therefore, the increase in net income under special order is calculated as $33,387.

(3)

Decrease in net income is calculated as follows:

Particulars
Amount
($93,000)
$33,387
($59,613)
Contribution lost
Net Income from Special order
Decrease in Net income

Therefore, decrease in net income is $59,613.

(Working notes :)

The contribution margin per unit is calculated as follows:

Contributionmarginperunit=SellingPriceVariablecost=Sellingprice(Directmaterials+Directlabor+Variablemanufacturingoverhead+Variablesellingexpenses)=$56($15+$8+$2+$1)=$30\begin{array}{c}\\{\rm{Contribution margin per unit = }}\,{\rm{Selling Price -- Variable cost}}\\\\{\rm{ = }}\,Selling\,price--\,\left( \begin{array}{l}\\Direct{\rm{ materials + Direct labor }}\\\\{\rm{ + Variable manufacturing overhead }}\\\\{\rm{ + Variable selling expenses}}\\\end{array} \right)\\\\ = \,\$ 56--\left( {\$ 15 + \$ 8 + \$ 2 + \$ 1} \right)\\\\ = \,\$ 30\\\end{array}

The contribution lost is calculated as follows:

Contributionlost=Unitsdropped×Contributionmarginperunit=3,100×$30=$93,000\begin{array}{c}\\{\rm{Contribution lost = Units dropped }} \times {\rm{ Contribution margin per unit}}\\\\{\rm{ = 3,100 }} \times {\rm{ \$ 30}}\\\\{\rm{ = }}\,\,{\rm{\$ 93,000}}\\\end{array}

Ans: Part 1

The total net income of the company due to this special order will increase by $73,098.

Part 2

The increase in net income under special order decision is $33,387.

Part 3

The decrease in net income under special order decision is $59,613.